Recently, Ms. James bought a house for $650,000. If the house market crashes and the value of the house decreases by 4% each year, determine what it will be worth in four years?

  1. Recently, Ms. James bought a house for $650,000. If the house market crashes and the value of the house decrease by 4% each year, determine what it will be worth in four years?
Answer #1

Worth after 4 years = $ 552075.26

We can use an exponential function to determine the value of house .

V(t) = Vo(1-r)t

Where,

Vo = Initial value 650,000

V(t) = Value of house at time t

r = decay rate , 4% = 0.04

t = time in years

Value after time t = 4 years

V(4) = 650000(1-0.04)4

= 650000(0.96)⁴

= 552075.26

Worth after 4 years = $ 552075.26


answered by: gavin